OpenStudy (anonymous):

Simplify when it is possible. 10x^2 -y + x^2 = 7x + 2x + 5y - 2y = 5x - 6y + 4x = 3ab + 4ab + 5 = 13a - 12b + a - b =

OpenStudy (whpalmer4):

Collect the like terms to simplify. For example, in the first one, you have two terms that are a number * x^2. Add the two numbers and multiply by x^2. 10x^2 - y + x^2 = 10x^2 - y + 1x^2 = (10+1)x^2 - y = 11x^2 - y